# USM RNN-T model weights binarization

### v1

@@ 1 Introduction
In the last several years the size of automatic speech recognition (ASR) models increased by more than 10x: from hundreds of millions [1, 2] to several billions parameters [3, 4, 5]. Serving cost goes up with model size increase, that is why ASR models compression is a hot research topic. Sparse network pruning [6, 7, 8, 9, 10] as well as quantization  [11, 12, 13, 14, 15] are successfully applied on end-to-end ASR. In [13, 14] authors show that it is possible to quantize 160M parameters model (trained on  900 hours of Librispeech data [16]) with 4 bits and 2 bits with no accuracy loss. But at the same time they show that a 160M parameters model quantized on a production data set (with millions of hours of speech) has 1.4% (absolute) WER degradation for 2 bits quantization [14]. Accuracy reduction can be even higher, e.g. in [17, 15] for 2 bits weights only quantization WER is increased by several times. To address this issue, authors in [15] explored the combination of quantization with sparsity. Based on these observation we can conclude that 4 bits weights quantization can be quality neutral on most of ASR models, 2 bits quantization works on some set of models (it depends on both model and data size and can require more experiments) and as expected 1 bit quantization (aka binarization) is the most challenging (e.g. authors of [17] reported  3x accuracy degradation), so it motivates us to investigate quantization aware weights binarization. Large speech models can be memory bound: memory communication latency can be comparable with computation latency. For example we observe 20% speed up (on TPU) due to 4 bit weights only quantization (quantize weights with int4 and then use float multiplication as in [13]). This is another reason to explore weights only binarization, so that we will be able to fit a larger model on a single TPU and get additional speed up due to reduction in memory communication.
@@ 5 Related work
As in [14] we explore quantization of RNN-T model (with 118M parameters) and train it on Librispeech data. We binarize weights of this model using best quantization method (from [14], based on absmax) and get only 0.1% WER increase on Librispeech test clean and 0.3% WER increase on test other accordingly. Note that we have to train this model two times longer in comparison to float baseline. As we can see in Table 2, binarization method Exp3 based on absmax, does not converge on a large model (USM RNN-T, with 1B parameters), trained on a large data set(5 million hours of speech). That is why in this work we are focused on large model binarization trained on large data.

### v2

@@ 1 Introduction
In the last several years the size of automatic speech recognition (ASR) models increased by more than 10x: from hundreds of millions [1, 2] to several billions parameters [3, 4, 5]. Serving cost goes up with model size increase, that is why ASR models compression is a hot research topic. Sparse network pruning [6, 7, 8, 9, 10] as well as quantization  [11, 12, 13, 14, 15] are successfully applied on end-to-end ASR. In [13, 14] authors show that it is possible to quantize 160M parameters model (trained on  900 hours of Librispeech data [16]) with 4 bits and 2 bits with no accuracy loss. But at the same time they show that a 160M parameters model quantized on a production data set (with millions of hours of speech) has 1.4% (absolute) WER degradation for 2 bits quantization [14]. Accuracy reduction can be even higher, e.g. in [17, 15] for 2 bits weights only quantization WER is increased by several times. To address this issue, authors in [15] explored the combination of quantization with sparsity. Based on these observation we can conclude that 4 bits weights quantization can be quality neutral on most of ASR models, 2 bits quantization works on some set of models (it depends on both model and data size and can require more experiments) and as expected 1 bit quantization (aka binarization) is the most challenging (e.g. authors of [17] reported  3x accuracy degradation), so it motivates us to investigate quantization aware weights binarization.
@@ 5 Related work
As in [14] we explore quantization of RNN-T model (with 118M parameters) and train it on Librispeech data. We binarize weights of this model using best quantization method (from [14], based on absmax) and get only 0.1% WER increase on Librispeech test clean and 0.3% WER increase on test other accordingly. Note that we have to train this model two times longer in comparison to float baseline. But as we can see in Table 2, binarization method Exp3 based on absmax, does not converge on a large model (USM RNN-T, with 1B parameters), trained on a large data set(5 million hours of speech). That is why in this work we are focused on large model binarization trained on large data.
@@ 6 Conclusion
7 Acknowledgements
We would like to thank Jian Li, Eugene Weinstein and Pedro Moreno for their support and valuable discussions.